OpenStudy (henrietepurina):

A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (henrietepurina):

@ParthKohli @Hero @undeadknight26

hero (hero):

Multiply the first equation by -3 then add it to the second equation.

OpenStudy (henrietepurina):

18x + -26y = 32 ???

hero (hero):

Not really

OpenStudy (henrietepurina):

Then what?

OpenStudy (henrietepurina):

@Hero

hero (hero):

What do you get if you multiplied both sides of the first equation by -3?

OpenStudy (henrietepurina):

x + (-21y) = 5.3 3x + (-1.6y) = -5.3

hero (hero):

Is this a new question you have posted?

OpenStudy (henrietepurina):

No, the same one

hero (hero):

If you don't know how to multiply, I'm not sure how much I'll be able to help you here.

OpenStudy (henrietepurina):

didn't i do it correctly?

hero (hero):

Let's start over...What is the first equation in the system?

OpenStudy (henrietepurina):

-3x + 7y = -16

hero (hero):

What result do you get after multiplying both sides of the first equation by -3?

OpenStudy (henrietepurina):

x + -21y = 48

hero (hero):

What happened to the coefficient of x?

OpenStudy (henrietepurina):

oh no, oops... 9 + -21y = 48

hero (hero):

Now the variable x is missing

hero (hero):

But anyway, you realize that you'll have 9x + -21y = 48 By the way, you should write these equations in their most simplest form: 9x - 21y = 48

hero (hero):

Anyway, take that result and add it to the second equation.

OpenStudy (henrietepurina):

ok.... - 26y = 32

OpenStudy (henrietepurina):

@hero amiright?

hero (hero):

Did you add or subtract? Looks to me like you added certain things but subtracted others.

OpenStudy (henrietepurina):

ohhh.... lemme try again

OpenStudy (henrietepurina):

-16y = 64

hero (hero):

That's better.

hero (hero):

So basically after you do that, you will be left with a system of -3x + 7y = -16 -16y = 64

hero (hero):

Now all you have to do is show that the system: -3x + 7y = -16 -9x + 5y = 16 Is equivalent to: -3x + 7y = -16 -16y = 64

OpenStudy (henrietepurina):

oh, so thats for 1?

OpenStudy (henrietepurina):

i meant a?

hero (hero):

If I have to answer that question, then it means you don't fully understand what I'm trying to explain to you.

OpenStudy (henrietepurina):

No, no I understand! Part A: -3x + 7y = -16 -16y = 64

hero (hero):

Yes, that's for part A. But that's how you should have asked if you're going to ask. You ask it this way: You say, "So is Part A: -3x + 7y = -16 -16y = 64 ?" And then, I will assume that you are only confirming what you already know.

OpenStudy (henrietepurina):

Yes :)

OpenStudy (henrietepurina):

Ok now im gonna get to part B...

hero (hero):

Okay, let me know how it goes

OpenStudy (henrietepurina):

-3x + 7y = -16 -9x + 5y = 16 9x + (-21y) = 48 -9x + 5y = 16 -21y) = 48 5y = 16 -16y = 64

OpenStudy (henrietepurina):

am i close @Hero

hero (hero):

That's the same thing we already did for Part A.

hero (hero):

In part B, what you should do is solve both systems using either elimination or substitution and see if you end up with the same (x,y) value for both systems.

OpenStudy (henrietepurina):

I dont get it?